Can someone help with this word problem: The profit on a watch is given by P = x2 – 13x – 80, where x is the number of watches sold per day. How many watches were sold on a day when there was a $50 loss?
Answer by stanbon(48558)

The profit on a watch is given by P = x^2 – 13x – 80, where x is the number of watches sold per day. How many watches were sold on a day when there was a $50 loss?
----------------
-50 = x^2-13x-80
x^2-13x-30=0
(x-3)(x-10)=0
x=3 or x=10
=============
Cheers,
Stan H.